OpenStudy (anonymous):

Patty and Mike are installing new floors in their house. Working alone, Mike can complete the floor in 10 hours. Patty can complete the same floor in 8 hours if working alone. How long will it take them, working together, to finish the floor? Round your answer to the nearest hundredth if necessary. Please help!!!

OpenStudy (anonymous):

\[\frac{ 1 }{ t }=\frac{ 1 }{ 10 }+\frac{ 1 }{ 8 }\] find t

OpenStudy (alurahenderson):

4.44 hours

OpenStudy (anonymous):

I got 40/9 for t

OpenStudy (anonymous):

alright yeah it's 4.44 thanks

OpenStudy (anonymous):

Can you help with another?

OpenStudy (alurahenderson):

sure but open up a new question